# Solving the Correlation Cluster LP in Sublinear Time

### v2

@@ 5 Finding One Small Ratio Cluster
$N_{\mathrm{cand}}(r)=\begin{cases}N_{\mathrm{adm}}(r)\setminus{V_{{\mathcal{K}}}}&\text{if }r\text{ is a singleton atom,}\\ K(r)\cup\left(\bigcap_{u\in K(r)}N_{\mathrm{adm}}(u)\right)&\text{if }r\text{ belongs to a non-singleton atom}.\end{cases}$
@@ 5.1 Overview of the Algorithm
where $\mathbb{1}$ is the indicator function for the event $v\in T$. Recall that we assume $uu$ is also a $+$edge in $G$. In the above calculation, the indicator variable corresponding to whether $v$ belongs to $T$ is used to adjust for the fact that the calculation differs slightly depending on whether $v$ is in $T$ or not. Since we do not know whether or not $v\in T$, we need this extra term.
@@
In this process, instead of sampling once, we sample $\eta$ different sets $A_{i}$. This is because we need new sampled nodes each time we add vertices to our final set $\hat{T}$. The algorithm also tries to guess the size of $C^{*}$. Since $C^{*}$ is an $\varepsilon$-large cluster, we know the size of $C^{*}$ will be within $[\varepsilon d^{+}(r),|N_{\mathrm{cand}}(r)|]$. We do not need the exact size, so we will enumerate possible sizes with different granularities, choosing values from the following set:
@@
11:        $T\leftarrow$ GenerateCluster($r,D(r),S^{1},\ldots,S^{\eta},\tilde{t}_{1},\ldots,\tilde{t}_{\eta}$)
@@
Algorithm 5 GenerateCluster($r,D(r),S^{1},\ldots,S^{\eta},\tilde{t}_{1},\ldots,\tilde{t}_{\eta}$)
1:  $T\leftarrow K(r),\hat{T}_{1}\leftarrow K(r)$
@@
$C^{*}_{i}=\hat{T}_{i}\cup\mathrm{argmin}\left\{\mathrm{cost}(\hat{T}_{i}\cup B)-w(\hat{T}_{i}\cup B)\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ B\subseteq\bigcup_{j=i}^{\eta}D_{r}^{j}\right\}.$
@@
GenerateCluster($r,D(r),S^{1},\ldots,S^{\eta},\tilde{t}_{1},\ldots,\tilde{t}_{\eta}$) produces a cluster $\hat{T}$ such that $\mathrm{cost}(\hat{T})\leq w(\hat{T})$.
@@
where $C$ is the $\varepsilon$-large cluster such that $K(r)\subseteq C\subseteq N_{\mathrm{cand}}(r)$. By Lemma 19, $|N_{\mathrm{cand}}(r)|=O(\varepsilon^{-3}d(r))$. If $r$ is in a non-singleton atom $K(r)\subseteq C^{\star}_{i}$, then $|C^{\star}_{i}|\geq|K(r)|\geq(1-O(\varepsilon))d(r)\geq\Omega(\varepsilon^{3}|N_{\mathrm{cand}}(r)|)$. $|K(r)|\geq(1-\varepsilon)d(r)$ is due to preclustering, every non-singleton atom has at most $O(\varepsilon)$ fraction $+$neighbors outside of $K(r)$.
@@
$-\frac{\gamma^{2}}{2}d_{\mathrm{adm}}(C)\geq\Delta(C^{\star}_{i})=\mathrm{cost}(C^{\star}_{i})-\sum_{v\in C^{\star}_{i}}w(v)\geq\sum_{v\in C^{\star}_{i}}d(v)-|C^{\star}_{i}|^{2}-\sum_{v\in C^{\star}_{i}}w(v).$
The last inequality uses $\mathrm{cost}(C^{\prime})+|C^{\prime}|^{2}\geq\sum_{v\in C^{\prime}}d(v)$, which holds for any cluster $C^{\prime}$. (The right side counts the $+$edges incident on $C^{\prime}$. Either such an edge leaves $C^{\prime}$ and contributes to the $\mathrm{cost}(C^{\prime})$ or it is inside $C^{\prime}$. There are at most $|C^{\prime}|^{2}/2$ edges inside $C^{\prime}$.) Since $C\subseteq N_{\mathrm{cand}}(r)$, $C$ only contains vertices in singleton atoms. Since $C$ is $\varepsilon$-large, all edges in $C$ are admissible. Thus, $d_{\mathrm{adm}}(C)\geq\frac{1}{2}|C|^{2}\geq\frac{\varepsilon^{2}}{2}d^{2}(r)$. Hence, $|C^{\star}_{i}|^{2}-\frac{\gamma^{2}}{2}d_{\mathrm{adm}}(C)\leq\varepsilon^{10}\gamma^{6}d^{2}(r)-\frac{\varepsilon^{2}\gamma^{2}}{4}d^{2}(r)\leq 0$. In particular,
$\sum_{v\in C^{\star}_{i}}d(v)-\sum_{v\in C^{\star}_{i}}w(v)\leq|C^{\star}_{i}|-\frac{\gamma^{2}}{2}d_{\mathrm{adm}}(C)\leq 0.$
There has to exist a vertex $v\in C^{\star}_{i}$ such that $d(v)-w(v)\leq 0$. ∎
@@
Lemma 21 outputs $\hat{T}$ such that $\mathrm{cost}(\hat{T})\leq w(\hat{T})$. However, we still need to satisfy the input conditions for Lemma 21.
@@ 5.2 Bounding $\Delta(C^{*}_{i+1})-\Delta(C^{*}_{i})$
$C^{*}_{i}=\hat{T}_{i}\cup\mathrm{argmin}\left\{\mathrm{cost}(\hat{T}_{i}\cup B)-w(\hat{T}_{i}\cup B)\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ B\subseteq\bigcup_{j=i}^{\eta}D_{r}^{i}\right\}.$
@@
${\mathbb{E}}[\leavevmode\nobreak\ \sum_{i=1}^{\eta_{0}}X_{i}\leavevmode\nobreak\ ]=\frac{\eta_{0}}{|T|}\cdot\frac{t}{\eta_{0}}\cdot d^{+}(v,T)=d^{+}(v,T).$
@@
$\Pr[\left|\leavevmode\nobreak\ \sum_{i=1}^{\eta_{0}}(X_{i}-{\mathbb{E}}[X_{i}])\leavevmode\nobreak\ \right|\geq\frac{\ell t}{\eta}]\leq 2\exp\left(-\frac{2(\ell t/\eta)^{2}}{\eta_{0}\cdot(t/\eta_{0})^{2}}\right)\leq 2\exp\left(-2\ell^{2}\eta\right).$
@@
$\quad\left|\mathrm{EstMarg}(S,\tilde{t},v)-\mathrm{Marginal}(T,v)\leavevmode\nobreak\ \right|$
$\leq\left|\tilde{t}-2\frac{d^{+}(v,S)}{|S|}\tilde{t}-|T|+2d^{+}(v,T)-\mathbb{1}(v\in T)\leavevmode\nobreak\ \right|$
$\leq\left|\tilde{t}-|T|\right|+2\frac{d^{+}(v,S)}{|S|}|\tilde{t}-t|+2\left|\leavevmode\nobreak\ \sum_{i=1}^{\eta_{0}}(X_{i}-{\mathbb{E}}[X_{i}])\leavevmode\nobreak\ \right|+1$
@@ 6 Refinements to Reach Nearly Linear Time
In this section, we present an algorithm to solve cluster LP in nearly linear time (i.e., $\widetilde{O}(m)$), where $m$ is the number of $+$edges. In a later section, we will present a sublinear-time algorithm. The main theorem regarding the nearly linear algorithm is stated as follows.
@@
Let $\varepsilon>0$ be a sufficiently small constant and let $\mathrm{cost}(\mathrm{OPT})$ be the cost of the optimum solution to the given Correlation Clustering instance. Then there is a small $\delta={\mathrm{poly}}(\varepsilon)$ such that the following statement holds: One can output a solution $(z_{S})_{S\subseteq V}$ to the cluster LP, described using a list of non-zero coordinates, with $\mathrm{obj}(x)\leq(1+\varepsilon)\mathrm{cost}(\mathrm{OPT})$ in expectation such that each coordinate of $z$ is either 00 or at least $\delta$. The running time to compute $z$ is $\widetilde{O}(2^{{\mathrm{poly}}(1/\varepsilon)}m)$.
@@
7:     If $\frac{{\mathrm{cover}}(K(v))}{p(K(v))}\leq(1+6\gamma)\leavevmode\nobreak\ R$, add $K(v)$ to $\mathcal{F}$ and set $\hat{p}_{w}=0$ for all $w\in K(v)$.
@@ 6.1 Approximate Ratio of Algorithm 6
Given vertex weights $p_{v}>0$, Algorithm 6 finds a family =$\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ such that,
@@
    for any distinct $S,T\in\mathcal{F}$, $S\cap T=\emptyset$,
@@
First, observe that Property 3. holds by the condition in the while loop. Property 2. holds since all small ratio clusters $\{C_{v}\}_{v\in U}$ do not split atoms. Next, we will prove the family $\mathcal{F}$ consists of disjoint sets. Observe that after we added a cluster $C\subseteq V$ to $\mathcal{F}$ we set the weight $\hat{p}_{v}$ to 00 for all $v\in C$. Note that if the weight $\hat{p}_{v}$ is set to 00 it remains 00 throughout the execution of the algorithm. Moreover, before we add a cluster $C\subseteq V$ to $\mathcal{F}$ we remove all vertices $v\in C$ with weight $p_{v}=0$ from $C$. Thus, when we add $C$ to $\mathcal{F}$ we have $\hat{p}_{v}>0$ for all $v\in C$ and $\hat{p}_{v}=0$ for all $v\in S,S\in\mathcal{F}$. Assume that the guess $R$ satisfies ${\mathrm{cover}}(\mathrm{OPT})\leq R\leq(1+\gamma){\mathrm{cover}}(\mathrm{OPT})$. It remains to prove the bound on the ratio $\frac{{\mathrm{cover}}(\mathcal{F})}{p(\mathcal{F})}$. To do so, we will make use of the following claim.
@@
Let $C^{\star}$ be the cluster among clusters in $\mathrm{OPT}$ with the smallest ratio $\frac{{\mathrm{cover}}(C^{\star})}{\hat{p}(C^{\star})}$. Notice that the algorithm will set some $\hat{p}_{v}$ to 00 without adding $u$ to $\mathcal{F}$, let
@@
For the second inequality, we use that $\gamma\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})\geq\gamma\leavevmode\nobreak\ \mathrm{cost}(\mathrm{OPT})\geq\gamma^{2}|E_{\text{adm}}|$ because of the preclustering (Theorem 7). The last inequality holds by the definition of $C^{\star}$. ∎
@@
$\frac{{\mathrm{cover}}(\mathcal{F})}{p(\mathcal{F})}=\frac{\sum_{S\in\mathcal{F}}{\mathrm{cover}}(S)}{\sum_{S\in\mathcal{F}}p(S)}\leq\frac{{\mathrm{cover}}(C_{v})}{\hat{p}(C_{v})}\leq(1+6\gamma)R\leq(1+8\gamma)\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})$
@@ 6.2 Runtime of Algorithm 6
$\frac{p^{(t)}_{u}}{4}\leq p^{(t+1)}_{u}\leq p^{(t)}_{u},\quad\text{if }u\text{is added to }\mathcal{F}.$
@@ 6.3 Wrap-Up: Proof of Nearly Linear Time Algorithm for cluster LP
We can obtain the preclustering $({\mathcal{K}},E_{\text{adm}})$ in time $\widetilde{O}(n)$ by Theorem 7. By Lemma 12, we can obtain a solution $z$ to the covering cluster LP after constant rounds $T_{\mathrm{MW}}=O({\mathrm{poly}}(1/\varepsilon))$ of Algorithm 2 such that ${\mathrm{cover}}(z)\leq(1+O(\gamma))\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})$. Moreover, $z_{S}$ is at least $\frac{1}{T_{\mathrm{MW}}}$ for each $S\in\mathrm{supp}(z)$ since each coordinate of the points $z^{(t)}$ is either 00 or at least $1$. The solution $z$ is simply the average of all the points $z^{(t)}$. Similar, the solution $z$ does not split atoms since each $z^{(t)}$ does not split atoms by Lemma 11. Furthermore, each vertex $v$ is contained in at most $T$ sets $S\in\mathrm{supp}(z)$ since the solutions $z^{(t)}$ have disjoint support by Lemma 11.
@@
$\mathrm{cost}(\widetilde{z})$ $\leq{\mathrm{cover}}(\widetilde{z})-d_{\mathrm{cross}}(V)$

### v5

@@ 5 Finding One Small Ratio Cluster
[…]
@@ 5.1 Overview of the Algorithm
where $\mathbb{1}$ is the indicator function for the event $v\in T$. Recall that we assume $uu$ is also a ++edge in $G$. In the above calculation, the indicator variable corresponding to whether $v$ belongs to $T$ is used to adjust for the fact that the calculation differs slightly depending on whether $v$ is in $T$ or not. Since we do not know whether or not $v\in T$, we need this extra term.
@@
In this process, instead of sampling once, we sample $\eta$ different sets $A_{i}$. This is because we need new sampled nodes each time we add vertices to our final set $\hat{T}$. The algorithm also tries to guess the size of $C^{*}$. Since $C^{*}$ is an $\varepsilon$-large cluster, we know the size of $C^{*}$ will be within $\varepsilon d^{+}(r),|N_{\mathrm{cand}}(r)|]$. We do not need the exact size, so we will enumerate possible sizes with different granularities, choosing values from the following set:
@@
[…]
@@
[…]
@@
$C^{*}_{i}=\hat{T}_{i}\cup\mathrm{argmin}\left\{\mathrm{cost}(\hat{T}_{i}\cup B)-w(\hat{T}_{i}\cup B)\penalty 10000\ \Big|\penalty 10000\ B\subseteq\bigcup_{j=i}^{\eta}D_{r}^{j}\right\}.$
@@
[…]
@@
where $C$ is the $\varepsilon$-large cluster such that $K(r)\subseteq C\subseteq N_{\mathrm{cand}}(r)$. By Lemma 19, $|N_{\mathrm{cand}}(r)|=O(\varepsilon^{-3}d(r))$. If $r$ is in a non-singleton atom $K(r)\subseteq C^{\star}_{i}$, then $|C^{\star}_{i}|\geq|K(r)|\geq(1-O(\varepsilon))d(r)\geq\Omega(\varepsilon^{3}|N_{\mathrm{cand}}(r)|)$. $|K(r)|\geq(1-\varepsilon)d(r)$ is due to preclustering, every non-singleton atom has at most $O(\varepsilon)$ fraction ++neighbors outside of $K(r)$.
@@
$\frac{\gamma^{2}}{2}d_{\mathrm{adm}}(C)\geq\Delta(C^{\star}_{i})=\mathrm{cost}(C^{\star}_{i})-\sum_{v\in C^{\star}_{i}}w(v)\geq\frac{1}{2}\sum_{v\in C^{\star}_{i}}d(v)-|C^{\star}_{i}|^{2}-\sum_{v\in C^{\star}_{i}}w(v).$
The last inequality uses $\mathrm{cost}(C^{\prime})+|C^{\prime}|^{2}\geq\frac{1}{2}\sum_{v\in C^{\prime}}d(v)$, which holds for any cluster $C^{\prime}$. (The right side counts the ++edges incident on $C^{\prime}$. Either such an edge leaves $C^{\prime}$ and contributes $1/2$ to the $\mathrm{cost}(C^{\prime})$ or it is inside $C^{\prime}$. There are at most $|C^{\prime}|^{2}/2$ edges inside $C^{\prime}$.) Since $\subseteq N_{\mathrm{cand}}(r), CC$ only contains vertices in singleton atoms. Since $C$ is $\varepsilon$-large, all edges in $C$ are admissible. Thus, $d_{\mathrm{adm}}(C)\geq\frac{1}{2}|C|^{2}\geq\frac{\varepsilon^{2}}{2}d^{2}(r)$. Hence, $|C^{\star}_{i}|^{2}-\frac{\gamma^{2}}{2}d_{\mathrm{adm}}(C)\leq\varepsilon^{10}\gamma^{6}d^{2}(r)-\frac{\varepsilon^{2}\gamma^{2}}{4}d^{2}(r)\leq 0$. In particular,
$\frac{1}{2}\sum_{v\in C^{\star}_{i}}d(v)-\sum_{v\in C^{\star}_{i}}w(v)\leq|C^{\star}_{i}|^{2}-\frac{\gamma^{2}}{2}d_{\mathrm{adm}}(C)\leq 0.$
There has to exist a vertex $v\in C^{\star}_{i}$ such that $\frac{1}{2}d(v)-w(v)\leq 0$. ∎
@@
[…]
@@ 5.2 Bounding $\Delta(C^{*}_{i+1})-\Delta(C^{*}_{i})$
$C^{*}_{i}=\hat{T}_{i}\cup\mathrm{argmin}\left\{\mathrm{cost}(\hat{T}_{i}\cup B)-w(\hat{T}_{i}\cup B)\penalty 10000\ \Big|\penalty 10000\ B\subseteq\bigcup_{j=i}^{\eta}D_{r}^{i}\right\}.$
@@
${\mathbb{E}}[\penalty 10000\ \sum_{i=1}^{\eta_{0}}X_{i}\penalty 10000\ ]=\frac{\eta_{0}}{|T|}\cdot\frac{t}{\eta_{0}}\cdot d^{+}(v,T)=d^{+}(v,T).$
@@
$\Pr[\left|\penalty 10000\ \sum_{i=1}^{\eta_{0}}(X_{i}-{\mathbb{E}}[X_{i}])\penalty 10000\ \right|\geq\frac{\ell t}{\eta}]\leq 2\exp\left(-\frac{2(\ell t/\eta)^{2}}{\eta_{0}\cdot(t/\eta_{0})^{2}}\right)\leq 2\exp\left(-2\ell^{2}\eta\right).$
@@
$\quad\left|\mathrm{EstMarg}(S,\tilde{t},v)-\mathrm{Marginal}(T,v)\penalty 10000\ \right|$
$\leq\left|\tilde{t}-2\frac{d^{+}(v,S)}{|S|}\tilde{t}-|T|+2d^{+}(v,T)-\mathbb{1}(v\in T)\penalty 10000\ \right|$
$\leq\left|\tilde{t}-|T|\right|+2\frac{d^{+}(v,S)}{|S|}|\tilde{t}-t|+2\left|\penalty 10000\ \sum_{i=1}^{\eta_{0}}(X_{i}-{\mathbb{E}}[X_{i}])\penalty 10000\ \right|+1$
@@ 6 Refinements to Reach Nearly Linear Time
In this section, we present an algorithm to solve cluster LP in nearly linear time (i.e., $\widetilde{O}(m)$), where $m$ is the number of ++edges. In a later section, we will present a sublinear-time algorithm. The main theorem regarding the nearly linear algorithm is stated as follows.
@@
Let $\varepsilon>0$ be a sufficiently small constant and let $\mathrm{cost}(\mathrm{OPT})$ be the cost of the optimum solution to the given Correlation Clustering instance. Then there is a small $\delta={\mathrm{poly}}(\varepsilon)$ such that the following statement holds: One can output a solution $(z_{S})_{S\subseteq V}$ to the cluster LP, described using a list of non-zero coordinates, with $\mathrm{obj}(x)\leq(1+\varepsilon)\mathrm{cost}(\mathrm{OPT})$ in expectation such that each coordinate of $z$ is either 0 or at least $\delta$. The running time to compute $z$ is $\widetilde{O}(2^{{\mathrm{poly}}(1/\varepsilon)}m)$.
@@
7:  If $\frac{{\mathrm{cover}}(K(v))}{p(K(v))}\leq(1+6\gamma)\penalty 10000\ R$, add $K(v)$ to $\mathcal{F}$ and set $\hat{p}_{w}=0$ for all $w\in K(v)$.
@@ 6.1 Approximate Ratio of Algorithm 6
Given vertex weights $p_{v}>0$, Algorithm 6 finds a family $\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ such that,
@@
    for any distinct $\in\mathcal{F}, ST=S\cap T$$\emptyset$,
@@
First, observe that Property 3. holds by the condition in the while loop. Property 2. holds since all small ratio clusters $\{C_{v}\}_{v\in U}$ do not split atoms. Next, we will prove the family $\mathcal{F}$ consists of disjoint sets. Observe that after we added a cluster $C\subseteq V$ to $\mathcal{F}$ we set the weight $\hat{p}_{v}$ to 0 for all $v\in C$. Note that if the weight $\hat{p}_{v}$ is set to 0 it remains 0 throughout the execution of the algorithm. Moreover, before we add a cluster $C\subseteq V$ to $\mathcal{F}$ we remove all vertices $v\in C$ with weight $p_{v}=0$ from $C$. Thus, when we add $C$ to $\mathcal{F}$ we have $\hat{p}_{v}>0$ for all $v\in C$ and $\hat{p}_{v}=0$ for all $v\in S,S\in\mathcal{F}$. Assume that the guess $R$ satisfies ${\mathrm{cover}}(\mathrm{OPT})\leq R\leq(1+\gamma){\mathrm{cover}}(\mathrm{OPT})$. It remains to prove the bound on the ratio $\frac{{\mathrm{cover}}(\mathcal{F})}{p(\mathcal{F})}$. To do so, we will make use of the following claim.
@@
Let $C^{\star}$ be the cluster among clusters in $\mathrm{OPT}$ with the smallest ratio $\frac{{\mathrm{cover}}(C^{\star})}{\hat{p}(C^{\star})}$. Notice that the algorithm will set some $\hat{p}_{v}$ to 0 without adding $u$ to $\mathcal{F}$, let
@@
For the second inequality, we use that $\gamma\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})\geq\gamma\penalty 10000\ \mathrm{cost}(\mathrm{OPT})\geq\gamma^{2}|E_{\text{adm}}|$ because of the preclustering (Theorem 7). The last inequality holds by the definition of $C^{\star}$. ∎
@@
$\frac{{\mathrm{cover}}(\mathcal{F})}{p(\mathcal{F})}=\frac{\sum_{S\in\mathcal{F}}{\mathrm{cover}}(S)}{\sum_{S\in\mathcal{F}}p(S)}\leq\frac{{\mathrm{cover}}(C_{v})}{\hat{p}(C_{v})}\leq(1+6\gamma)R\leq(1+8\gamma)\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})$
@@ 6.2 Runtime of Algorithm 6
[…]
@@ 6.3 Wrap-Up: Proof of Nearly Linear Time Algorithm for cluster LP
We can obtain the preclustering $({\mathcal{K}},E_{\text{adm}})$ in time $\widetilde{O}(n)$ by Theorem 7. By Lemma 12, we can obtain a solution $z$ to the covering cluster LP after constant rounds $T_{\mathrm{MW}}=O({\mathrm{poly}}(1/\varepsilon))$ of Algorithm 2 such that ${\mathrm{cover}}(z)\leq(1+O(\gamma))\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})$. Moreover, $z_{S}$ is at least $\frac{1}{T_{\mathrm{MW}}}$ for each $S\in\mathrm{supp}(z)$ since each coordinate of the points $z^{(t)}$ is either 0 or at least $1$. The solution $z$ is simply the average of all the points $z^{(t)}$. Similar, the solution $z$ does not split atoms since each $z^{(t)}$ does not split atoms by Lemma 11. Furthermore, each vertex $v$ is contained in at most $T$ sets $S\in\mathrm{supp}(z)$ since the solutions $z^{(t)}$ have disjoint support by Lemma 11.
@@
[…]